3. Find the general solution of the partial differential equation

6uxx −5uxy −6uyy =52e^3x+2y
1
Expert's answer
2020-04-14T13:43:46-0400

1 STEP: We rewrite this equation using the notation



"\frac{\partial}{\partial x}\equiv D\quad\text{and}\quad\frac{\partial}{\partial y}=D'\longrightarrow\\[0.3cm]
6u_{xx}-5u_{xy}-6u_{yy}=52\cdot e^{3x+2y}\longrightarrow\\[0.3cm]
\boxed{\left(6D^2-5DD'-6D'^2\right)u=52\cdot e^{3x+2y}}"

2 STEP: Factor the resulting square expression

Note: I will use the letters "x" and "y" so that there is no confusion between the operators "D" , "D'" and the designation of the discriminant "D" :



"6D^2-5DD'-6D'^2\longrightarrow 6x^2-5yx-6y^2=0\\[0.3cm]
\left.\begin{array}{l}
a=6\\
b=-5y\\
c=-6y^2
\end{array}\right| D=b^2-4ac=(-5y)^2-4\cdot 6\cdot(-6y)\\[0.3cm]
D=25y^2+144y^2=169y^2\longrightarrow\sqrt{D}=13y\\[0.3cm]
\left[\begin{array}{l}
x_1=\displaystyle\frac{-b-\sqrt{D}}{2a}=\displaystyle\frac{5y-13y}{12}=-\displaystyle\frac{2y}{3}\\[0.3cm]
x_2=\displaystyle\frac{-b+\sqrt{D}}{2a}=\displaystyle\frac{5y+13y}{12}=\displaystyle\frac{3y}{2}
\end{array}\right.\\[0.3cm]
6x^2-5xy-6y=6\cdot\left(x-\left(-\frac{2y}{3}\right)\right)\cdot\left(x-\frac{3y}{2}\right)\\[0.3cm]
\boxed{6D^2-5DD'-6D'^2=6\cdot\left(D+\frac{2}{3}\cdot D'\right)\cdot\left(D-\frac{3}{2}\cdot D'\right)}"

3 STEP: A bit of theory (without proof)

Suppose we need to solve a partial differential equation



"f\left(D,D'\right)u=\varphi(x,y)"

The solution consists of two parts:



"u(x,y)=C.F.+P.I.,\,\,\text{where}\\[0.31cm]
\left[\begin{array}{l}
C.F.-\text{the complementary function}\\[0.3cm]
P.I.-\text{Particular Integral }
\end{array}\right."



To find the complementary function (C.F.)

Put "D=m" and "D'=1" in "f\left(D, D'\right)=0" , then we get an equation, which is called the auxillary equation.

∴ the auxillary equation is "f(m,1)=0" i.e.,



"a_0m^n + a_1m^{n\u22121} + \u00b7 \u00b7 \u00b7 + a_n = 0"

Let the roots of this equation be "m_1" , "m_2" , . . . , "m_n" . If the roots are real (or imaginary) and distinct, then



"C.F.=f_1(y+m_1x)+f_2(y+m_2x)+ \u00b7 \u00b7 \u00b7 +f_n(y+m_nx)"



To find Particular Integral (P.I.)

If



"\varphi(x, y)=e^{ax+by}"

then



"P.I.=\frac{1}{f\left(D,D'\right)}\cdot e^{ax+by}=\frac{1}{f(a,b)}\cdot e^{ax+by},\text{provided} f(a, b)\neq0."

If "f(a,b)=0" , then


"P.I.=x\cdot\frac{1}{f'\left(D, D'\right)}\cdot e^{ax+by},\\[0.3cm]
\text{where}\,\,\,f'\left(D, D'\right)\,\,\text{is the partial derivative of} \,\,f\left(D, D'\right) w.r.t \,\,D."

4 STEP: Put theory into practice


"С.F. :\left(D+\frac{2}{3}\cdot D'\right)\cdot\left(D-\frac{3}{2}\cdot D'\right)u=0\longrightarrow\\[0.3cm]
\boxed{C.F.=f_1\left(y-\frac{2x}{3}\right)+f_2\left(y+\frac{3x}{2}\right)}\\[0.3cm]
P.I. : P.I.=\frac{1}{\left(D+\frac{2}{3}\cdot D'\right)\cdot\left(D-\frac{3}{2}\cdot D'\right)}\cdot\frac{52}{6}\cdot e^{3x+2y}\\[0.3cm]
P.I.=\frac{1}{\left(3+\frac{2}{3}\cdot 2\right)\cdot\left(3-\frac{3}{2}\cdot 2\right)}\cdot\frac{52}{6}\cdot e^{3x+2y}\\[0.3cm]
P.I.=\frac{1}{\left(3+\frac{4}{3}\right)\cdot\underbrace{\left(3-3\right)}}_{=0}\cdot\frac{52}{6}\cdot e^{3x+2y}\\[0.3cm]"

This means that you must first differentiate the expression "\left(6D^2-5DD'-6D'^2\right)" by "D" , so



"\frac{\partial}{\partial D}\left(6D^2-5DD'-6D'^2\right)=12D-5D'"

Then,



"P.I. : P.I.=\frac{1}{12D-5D'}\cdot 52\cdot e^{3x+2y}\\[0.3cm]
P.I.=\frac{1}{12\cdot3-5\cdot 2}\cdot 52\cdot e^{3x+2y}\\[0.3cm]
P.I.=\frac{52\cdot e^{3x+2y}}{26}\equiv 2\cdot e^{3x+2y}\\[0.3cm]"

Conclusion,



"u(x,y)=C.F.+P.I.\longrightarrow\\[0.3cm]
\boxed{u(x,y)=f_1\left(y-\frac{2x}{3}\right)+f_2\left(y+\frac{3x}{2}\right)+2\cdot e^{3x+2y}}"



ANSWER



"u(x,y)=f_1\left(y-\frac{2x}{3}\right)+f_2\left(y+\frac{3x}{2}\right)+2\cdot e^{3x+2y}"